Opened 9 years ago

Closed 2 years ago

#386 closed flight processing (fixed)

EM10/07, flight day 145b/2010, Trapani

Reported by: anch Owned by: anch
Priority: alpha 4 low Milestone: 2010 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by knpa)

Specific Scientific Objectives

  1. Development and validation of hyperspectral retrieval algorithms for biophysical parameter estimation in preparation for launch of PRISMA hyperspectral mission. These include;

Leaf area indexn (LAI m2/m2)
Fraction of vegetation cover (Fc, %)
Leaf chlorophyll concentration and canopy chlorophyll content.

  1. Test and validation of remote sensing techniques for qualifying water exchanges in the main tree crop species of the Med.

PRIORITY:
a4l

Data location: /users/rsg/arsf/workspace/EM10_07-2010_145a_Trapani

Data arrived from ARSF via network transfer on 08/06/2010

PI: Dr F. Vuolo

Sensors:

  • Eagle (16/02/2011)
  • Hawk (16/02/2011)
  • Leica LIDAR (04/01/12)
  • RCD Photography (Not requested)

Attachments (1)

1452010.sync (354 bytes) - added by anch 9 years ago.

Download all attachments as: .zip

Change History (30)

comment:1 Changed 9 years ago by anch

  • Summary changed from EM10/07, flight day 143b/2010, Trapani to EM10/07, flight day 145b/2010, Trapani

comment:2 Changed 9 years ago by anch

Starting navigation processing

comment:3 Changed 9 years ago by anch

Applanix and IPAS processing completed.

Changed 9 years ago by anch

comment:4 Changed 9 years ago by anch

Hyperspectral complete and delivery created. Submitted for delivery check

comment:5 Changed 9 years ago by emca

Hyperspectral delivery check complete.

All data looks good - just need to change the flight letter from 'a' to 'b' on the screenshots, otherwise ready to deliver.

comment:6 Changed 9 years ago by anch

  • Description modified (diff)

Delivered Eagle and Hawk.

Sent to Dr. Vuolo in Austria and Dr. Dash in Southampton

comment:7 Changed 8 years ago by knpa

Starting LiDAR processing.

comment:8 Changed 8 years ago by knpa

Just noticed that this project has RCD data which is not recorded on the status page.
Can not add it right now due to bug, this will need to be done, and the RCD processed, before this project is archived.

comment:9 Changed 8 years ago by knpa

Bug has been fixed (thanks mggr), RCD added to status page.

comment:10 Changed 8 years ago by knpa

145b is actually two sites, 9km apart.

Will keep together as this is how it was previously delivered (HS).

comment:11 Changed 8 years ago by knpa

All lines look good and line up well with each other.

comment:12 Changed 8 years ago by knpa

QC completed.

Preparing delivery.

comment:13 Changed 8 years ago by knpa

LiDAR ready for delivery check at: ~airborne/workspace/EM10_07-2010_145b_Trapani/delivery/20111230

comment:14 Changed 8 years ago by adbe

Delivery Check:

  • Principal Investigator: Francesco Vuolo instead of Dr F. Vuolo in ReadMe
  • DEM is 5 metre res(?), ReadMe says 2 metre

Otherwise ready to deliver

comment:15 Changed 8 years ago by knpa

LiDAR data delivered today to Francesco Vuolo at BOKU in Vienna (04/01/12)

comment:16 Changed 8 years ago by knpa

  • Description modified (diff)
  • Milestone set to 2010 data processing completion

comment:17 Changed 8 years ago by jaho

Processing RCD.

comment:18 Changed 8 years ago by jaho

  • Moved camera pictures from 145a to 145b as they seem to belong here.
  • It appears that GPS times in ImageEvents file are given in day seconds rather than week seconds. Need to convert before processing.
  • Removed camera images 1 - 14, no corresponding flightlines.

comment:19 Changed 8 years ago by knpa

Starting reprocessing.

comment:20 Changed 8 years ago by knpa

Stage 1 complete.

Problem with stage 2 - apl claiming that it can't find a navigation file to match the eagle raw files.

comment:21 Changed 8 years ago by knpa

Found the problem - there is only one eagle nav file for this day and it is living in 145a.

Symlinked and running again with hawk as well as the bad pixel stuff is finished now.

comment:22 Changed 8 years ago by knpa

Lev3's look OK.

Running stage 3.

comment:23 Changed 8 years ago by knpa

Realised that no line information files had been created for this project.

This is due to an issue with the database, where there is only an entry for 145 and not 145b.

Have created manually using the same command with the --force option.

comment:24 Changed 7 years ago by knpa

Deleting project and re-running stage 2 as apl chain has been updated to generate additional masked files.

comment:25 Changed 7 years ago by knpa

Re-ran stage 2 for Hawk as aplcal has been updated in relation to the bad pixel file.

comment:26 Changed 7 years ago by emca

Delivery checking reprocessed apl data.

  • DEM should be named EM10_07-2010_145b-ASTER.hdr
  • Logsheet is joint for a + b
  • Flightline names in logsheet don't match readme (cross lines)

Zipping mapped files. Waiting for check_apl_command to finish.

comment:27 Changed 7 years ago by emca

Checked apl commands, data looks fine. Delivery check complete.

comment:28 Changed 6 years ago by knpa

Merged and deleted workspace version.

comment:29 Changed 2 years ago by asm

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Archived and it is available from CEDA. Re-processed hyperspectral data at http://browse.ceda.ac.uk/browse/neodc/arsf/2010/EM10_07/EM10_07-2010_145b_Trapani/hyperspectral-reprocessed

Plymouth 12th May 2017

Note: See TracTickets for help on using tickets.